An ideal gas initially having volume 2.0 x 10-3m3 at a pressure of 1.2 x 105 Pa and a temperature of 0℃ is brought around a cycle. (A) The pressure is increased to 1.5 x 105 Pa without a change in volume, (B) the gas expands isothermally until it returns to its original pressure and (C) the gas is compressed at constant pressure until it returns to its original state.


What is (a) the temperature of the gas during process B, (b) the volume of gas at the end of process B and (c) the work done during the cycle.


1
Expert's answer
2021-04-23T07:27:13-0400

Solution.

"V_1=2.0\sdot10^{-3}m^3;"

"p_1=1.2\sdot10^5Pa;"

"T_1=273K;"

"p_2=1.5\sdot10^5Pa;"

"a) V=const;"

"\dfrac{P_1}{T_1}=\dfrac{P_2}{T_2}\implies T_2=\dfrac{P_2\sdot T_1}{P_1};"

"T_2=\dfrac{1.5\sdot10^5Pa\sdot 273K}{1.2\sdot10^5Pa}=341K;"

"b) P_1V_1=P_2V_2\implies V_2=\dfrac{P_1V_1}{P_2};"

"V_2=\dfrac{1.5\sdot10^5Pa\sdot2\sdot10^{-3}m^3}{1.2\sdot10^5Pa}=2.5\sdot10^{-3}m^3;"

"c)" "A=\nu RT_2ln\dfrac{V_2}{V_1}-p_1\Delta V;"

"\nu=\dfrac{P_1V_1}{T_1};"

"\nu=\dfrac{1.2\sdot10^5Pa\sdot 2.0\sdot10^{-3}m^3}{273K}=0.88mol;"

"A=0.88mol\sdot8.31J/(molK)441Kln\dfrac{2.5\sdot10^{-3}m^{-3}}{2\sdot10^{-3}m^{-3}}+1.2\sdot10^5\sdot0.5m^{-3}=-1635J;"

Answer: "a)T_2=341K;"

"b)" "V_2=2.5\sdot10^{-3}m^3;"

"c)A=-1635J."